# Does Data Contamination Detection Work (Well) for LLMs? A Survey and Evaluation on Detection Assumptions

### v1

@@ 5 Other Types of Contamination
Besides direct data contamination, previous research also investigates indirect data contamination (6 papers), and task contamination (5 papers).
@@ 5.1 Indirect Data Contamination
Indirect data contamination occurs when an instance $x$ is similar to an instance $x^{\prime}\in D_{M}$, but with a relatively lower degree of similarity compared to direct contamination. For instance, $x$ might be a paraphrase or summary of $x^{\prime}$ Yang et al. (2023).
In practice, indirect data contamination often originates from sources that are not explicitly recognized as part of the original dataset complicating the tracking and tracing process Balloccu et al. (2024). For example, OpenAI uses online user conversations for training, which could include variations of benchmark datasets  [^5] https://help.openai.com/en/articles/5722486-how-your-data-is-used-to-improve-model-performance. Another example involves knowledge distillation, where an LM utilizes instances $x_{k}$ generated by another LM $M^{\prime}$ during training, and these instances $x_{k}$ may resemble instances from the training set $D_{M^{\prime}}$ of $M^{\prime}$ Veselovsky et al. (2023).
@@ 5.1.1 Detection Approaches for Indirect Data Contamination
However, three approaches may still be applicable: (1) the instance similarity measured by another LM Yang et al. (2023), (2) the CDD metric proposed by Dong et al. (2024), which leverages Assumption A6 by measuring output variations rather than directly comparing with original instances, and (3) directly tracking the disclosed usage of datasets. For example, Balloccu et al. (2024) reviewed the datasets evaluated using OpenAI APIs.
@@ 5.2 Task Contamination
Task contamination occurs when any instance of the same task is seen by an LM Li and Flanigan (2024). Detecting task contamination is crucial for assessing an LM’s generalizability to unseen tasks Chung et al. (2024). Tasks can include applications such as machine translation, summarization, and mathematical calculation. Task contamination is a broader concept than data contamination: if a dataset is partially contaminated, the associated task is contaminated, but task contamination doesn’t always imply the dataset has been seen.
@@
For example, Ranaldi et al. (2024) and Li and Flanigan (2024) find that OpenAI models perform significantly better on benchmarks released before the model’s release than on those released later. To ensure fair comparisons across tasks, Li and Flanigan (2024) control task difficulty using a baseline system. However, Cao et al. (2024) also note that LMs do not necessarily perform worse on more recent code generation benchmarks.
As a variation of this assumption, Aiyappa et al. (2023) find improved performance on the same benchmark with model updates.
@@ 6 Case Study
Besides the case study in Section 4.1, we aim to evaluate if the assumptions outlined in Table 1 are universally applicable across different domains, for direct and instance-level data contamination.
@@
Github
Free-
Law
ArXiv
PubMed
Central
Enron
Emails
PubMed
Abstracts
Open-
Subtitles
OpenWeb-
Text2
@@
Hacker-
News
Pile-CC
A1 PPL_50 49.7 48.9 50.7 49.4 49.6 49.7 47.9 50.3 50.8 48.4 49.8
PPL_100 50.8 49.4 51.1 50.1 50.5 48.8 48.1 48.9 50.3 47.3 50.4
PPL_200 51.1 49.4 50.9 50.2 51.4 49.9 47.8 48.5 49.5 50.7 52.1
Min 5% token 51.5 49.8 51.3 48.5 50.0 48.5 49.8 49.7 49.0 47.8 50.9
Min 15% token 51.2 49.8 51.3 49.2 50.1 49.4 49.9 49.2 49.2 49.6 51.1
Min 25% token 51.0 49.4 51.4 49.5 50.0 49.6 50.2 49.0 49.2 49.8 51.1
A4 Mem 1 49.2 48.9 50.5 47.1 49.8 50.7 49.6 50.6 49.4 48.3 49.7
Mem 3 49.7 50.0 51.0 46.5 50.1 49.8 51.5 50.5 48.2 49.3 48.7
Mem 5 49.7 48.0 50.8 48.2 49.7 52.7 50.8 50.2 49.1 50.0 49.4
A6 Entropy 5 51.9 49.2 52.4 48.9 50.3 48.0 49.4 51.9 48.6 49.0 50.4
Entropy 10 52.2 49.4 52.3 49.3 50.4 48.2 48.7 52.1 48.6 49.0 50.6
Entropy 25 52.1 49.5 50.9 49.5 50.7 48.5 48.1 51.0 48.1 49.6 50.9
PPL Seen 3.0$\pm$5.6 3.9$\pm$1.4 5.6$\pm$2.4 5.6$\pm$2.1 7.8$\pm$10.6 8.7$\pm$4.4 9.5$\pm$3.4 10.8$\pm$9.2 11.0$\pm$13.2 12.1$\pm$5.9 13.9$\pm$11.8
Unseen 3.0$\pm$6.2 4.0$\pm$3.0 5.7$\pm$2.1 5.6$\pm$2.2 8.0$\pm$10.4 8.6$\pm$4.3 9.6$\pm$3.6 10.6$\pm$8.2 10.8$\pm$13.2 11.8$\pm$5.5 13.9$\pm$10.5
Table 2: Average contamination detection AUC for the Pythia-6.9b model, under different PILE domains. ‘PPL’ represents the average perplexity $\pm$ STD, where instances are truncated by the model max length of 2048 tokens. The columns (domains) are sorted by the average PPL among the seen instances.
@@ 6.1 Assumptions to Evaluate
Because some assumptions have specific requirements and their applicability depends on how well these requirements are met, we select the three assumptions without any requirements to evaluate. Thus, we restrict the confounding factors and leave the testing of other assumptions for future research.
Assumption A1: Absolute Probability. In the assumption that seen instances will have a lower perplexity and higher Min Top $p$ Probability, we measure perplexity by an instance’s first $k$ tokens (PPL_$k$) Carlini et al. (2021); Min Top $p$ Prob by the average of the least likely tokens, whose cumulative probabilities are $p\%$ Shi et al. (2023).
Assumption A4: Verbatim Memorization. We expand this assumption from the instance level to the token level, assuming seen instances will memorize more tokens. We measure the percentage of tokens in an instance ranked as the $k$ most likely in casual language modeling (Mem $k$). The $k$ value of 1 represents greedy decoding, and larger than 1 simulates the decoding with top $k$ token sampling.
Assumption A6: Generation Variation. We evaluate the related assumption to generation variation that seen instances will have better certainty and lower entropy. We measure entropy over the top $k$ most likely tokens (Entropy $k$) (see Appendix A.4 for details).
@@ 6.2 Experiment Design
We utilize the widely used Pythia LMs of different sizes [^6] https://huggingface.co/EleutherAI/pythia-70m Biderman et al. (2023), which are pretrained on the 825GiB Pile dataset [^7] https://Pile.eleuther.ai/, https://huggingface.co/datasets/ArmelR/the-Pile-splitted Gao et al. (2020).
Because the Pile is composed of 22 smaller subsets from different domains, we randomly sample 11 domains for our experiment.Within each domain, we randomly select 1k instances from the train split (seen), and 1k instances from the test split (unseen). We evaluate the detection performance through the area under the curve (AUC), at the instance level.
@@ 6.3 Results
6.3.1 Within-Domain Detection
Table 2 shows the AUC for each contamination detection method on Pythia-6.9b. Within the same domain, the similar average PPL between seen and unseen instances indicates that they have similar underlying distributions, but also a high variation (STD). However, PPL differs a lot across domains.
We also observed all metrics perform close to random guessing, with AUC close to 50. Our results for Assumption A1 are consistent with critiques they received (see Section 4.2.1). We conclude that the LM learns underlying data distributions rather than memorizing specific instances. Among different sizes of Pythia models, we find that larger LMs have a lower average PPL in the same domain. However, all metrics have similar, near-random performance (see Appendix A.5.1).
6.3.2 Cross-Domains Detection
As Table 2 suggests domain difference greatly impacts PPL, we further evaluate the scenario where seen and unseen instances are from different domains, with the same instances in Section 6.2.
Figure 1: Average contamination detection AUC for the Pythia-6.9b model with PPL_200, when the seen and unseen instances are from different domains.
We present the AUC with PPL_200 in Figure 1. The AUC is high in the top-right corner when seen instances have high average PPL and unseen instances have low average PPL. Conversely, the bottom-left corner has low AUCs. This indicates that the PPL_200 detects domain shifts, instead of data contamination. We additionally visualize the PPL_200 distribution within and across domains in Appendix A.5.2. A similar trend is observed with other metrics (see Appendix A.5.2 & A.5.3).
@@ 7 Conclusion
In this study, we present a comprehensive survey of 47 studies focused on data contamination detection and their underlying assumptions. Our analysis reveals that these assumptions may not apply consistently across different contexts. Our case studies confirmed that 3 out of the 8 assumptions are not universally applicable across all training domains.
We conclude that many assumptions measure an LM’s goodness of fit, which is not necessarily the result of instance memorization due to data contamination. Therefore, detecting direct data contamination remains challenging.
More generally, this study highlights the importance of researchers clearly stating and validating the assumptions behind their proposed approaches, because approaches built on shaky assumptions are unlikely to perform well.
@@ 8 Limitations
In this study, we reviewed 47 papers on data contamination detection, but there may be additional relevant studies not captured by our collection methods. We primarily focus on data contamination detection approaches for English LMs. Other detection approaches for non-English LMs, data modalities beyond text, and other machine learning techniques may exist and could potentially be transferable to English LMs.
Our case study deployed the Pythia model family, which is trained for no longer than 2 epochs for each domain, and has therefore seen each instance for at most twice. In this case, although an instance is contaminated, it is difficult for the LM to memorize the instance. Those assumptions might perform very differently for other LMs and other datasets.
@@ 9 Ethical Considerations
In this work, we employed the Pythia model family Biderman et al. (2023) and their pretraining corpus, the 825 GiB Pile dataset Gao et al. (2020). Since the Pile includes content crawled from the Internet, it may contain sensitive and/or identifiable information. Therefore, we only downloaded the necessary instances and published the numerical results from our experiments. Upon acceptance of our work, we will release our code for data sampling and contamination detection approaches to ensure reproducibility, but we will not publish any actual data instances or language models. We advise the community to check data regulations before deploying the Pythia model family Biderman et al. (2023) and the Pile dataset Gao et al. (2020) for other purposes.
@@ A.1 Table of Notations
$x$ An language instance, as a series of tokens.
@@
contamination, which takes the input as two
instances $x$ and $x^{\prime}$ and outputs False (0) or
@@
two instances, $x$ and $x^{\prime}$ and outputs and
outputs a real value.
@@
contamination, defined by Equation 1.
@@
least likely tokens in an instance $x$, for
a given LM $M$.
$\tau_{\cdot}$
@@
for $M$ under diverse, different sampling
strategies. given $x_{p}$
@@
The entropy of the top $k$ most likely tokens,
defined by Equation 11.
@@ A.2 Risks and Mitigation Approaches for Data Contamination
During our paper collection, we identify the relevant research on the risks and mitigation strategies for data contamination, which does not involve proposing or evaluating existing detection approaches for data contamination. While excluding those papers from the main text of this manuscript, we present them in Table 4 in this Appendix.

### v3

@@ 5 Other Types of Contamination
Besides direct data contamination, previous research also investigates indirect data contamination (6 papers) and task contamination (5 papers).
@@ 5.1 Indirect Data Contamination
Indirect data contamination occurs when an instance $x$ is not seen by an LM $M$, but something ($x^{\prime}$) derived from x is. For instance, $x^{\prime}$ can be a paraphrase or a summary of $x$ Yang et al. (2023).
Indirect data contamination is often hard to track and trace Balloccu et al. (2024). For example, OpenAI uses online user conversations for training, which could include variations of benchmark datasets [^5] https://help.openai.com/en/articles/5722486-how-your-data-is-used-to-improve-model-performance. Accessed on Oct 6, 2024.. Another example involves knowledge distillation, where an LM utilizes instances $x_{k}$ generated by another LM $M^{\prime}$ during training, and these instances $x_{k}$ may resemble instances from the training set $D_{M^{\prime}}$ of $M^{\prime}$ Veselovsky et al. (2023).
@@ 5.1.1 Detection Approaches for Indirect Data Contamination
However, prior research showed that three approaches may still be applicable: (1) the instance similarity measured by another LM Yang et al. (2023), (2) the CDD metric Dong et al. (2024), which leverages Assumption A6 by measuring output variations rather than directly comparing with original instances, and (3) directly tracking the disclosed usage of datasets. For example, Balloccu et al. (2024) reviewed the datasets evaluated using OpenAI APIs.
@@ 5.2 Task Contamination
Task contamination occurs when any instance of the same task is seen by an LM Li and Flanigan (2024). Detecting task contamination is crucial for assessing an LM’s generalizability to unseen tasks Chung et al. (2024). Tasks can include applications such as machine translation, summarization, and mathematical calculation. Task contamination is a broader concept than data contamination: if some labeled instances from a dataset are seen by an LM, the associated task is contaminated, but task contamination doesn’t always imply the dataset has been seen.
@@
For example, as noted by Aiyappa et al. (2023), LLMs show improved performance on the same benchmark after model updates, which may be influenced by data contamination during LLMs’ continuous training. Ranaldi et al. (2024) and Li and Flanigan (2024) also find that OpenAI models perform significantly better on benchmarks released before the model’s release than on those released later, when task difficulty is controlled or performance is normalized using a baseline model. To ensure fair comparisons across tasks, Li and Flanigan (2024) control task difficulty using a baseline system. However, Cao et al. (2024) also note that LMs do not necessarily perform worse on more recent code generation benchmarks.
@@ 6 Case Study
Besides the case study in Section 4.1, we aim to evaluate whether the assumptions outlined in Table 1 are universally applicable across different domains, for direct and instance-level MIA.
LM
Params
Training
Phase
Epochs
Trainset
Size
Batch
Size
Seen & Unseen Datasets
Used in Our Case Study
Pythia Biderman et al. (2023) 70M - 12B Pretraining $\approx$ 1.5 825 GiB 2M Pile Gao et al. (2020)
OLMo-2 OLMo et al. (2024) 7B $\approx$ 2 22.4 TB $\approx$ 4M AlgebraicStack Azerbayev et al. (2023)
Zephyr-7B-$\beta$ Tunstall et al. (2023) 7B Supervised Fine-tuning 1-3 9.3 GB 512 UltraChat Ding et al. (2023)
BioMistral-NLU Fu et al. (2025) 7B 2 3.6 GB 64 Medical-NLU Fu et al. (2025)
Table 2: LMs and datasets used in the case study. Except for the UltraChat dataset, each dataset contains multiple subsets from different domains. The trainset refers to the whole trainset used in each LM’s corresponding training phase, as described in their original paper, which is a superset of seen & unseen datasets used in our case study.
Training Phase Pretaining Supervised Fine-tuning
Model Pythia-6.9B OLMO-2-7B Zephyr-7B-$\beta$ BioMistral-NLU-7B
@@
ArXiv
Github-
Coq
Github-
Isabelle
Ultra
Chat
DC-
MTSample
RE-
2012temp
PPL_50 50.7 50.7 47.1 50.9 55.3 51.9 62.5
PPL_100 50.4 51.1 48.2 53.1 59.1 60.0 95.3
PPL_200 49.6 50.9 48.5 51.6 60.1 58.9 99.4
Min 5% token 48.5 51.3 49.4 54.0 63.6 47.4 92.9
Min 15% token 48.6 51.3 49.3 53.6 63.0 51.7 93.3
A1 Min 25% token 48.5 51.4 49.3 53.1 61.5 53.3 93.4
Mem 5 49.1 52.7 52.1 48.7 52.2 47.9 41.4
Mem 15 48.6 51.9 50.6 58.6 53.1 49.1 45.2
A4 Mem 25 48.2 51.6 49.4 59.2 53.1 50.3 48.9
Entropy 5 49.3 52.0 47.2 52.3 54.3 55.4 93.2
Entropy 15 49.0 52.0 47.8 52.1 54.1 54.8 93.1
A6 Entropy 25 48.9 51.9 48.6 52.5 54.0 54.1 93.1
Average AUC 49.0 51.8 49.0 53.3 55.9 52.0 74.1
Seen 13.2$\pm$16.0 7.9$\pm$3.7 10.5$\pm$8.1 8.4$\pm$5.2 5.3$\pm$4.6 1.7$\pm$0.2 1.4$\pm$0.1
PPL Unseen 12.7$\pm$10.6 8.0$\pm$3.6 9.9$\pm$7.2 9.2$\pm$6.5 6.2$\pm$4.1 1.8$\pm$0.2 3.0$\pm$1.6
Table 3: Average MIA AUC for different LMs. For LMs evaluated on multiple subsets (domains) of the same dataset, we present the results from the subsets with the lowest and highest average AUC. The last two rows, marked as ‘PPL_200’, represent the average perplexity $\pm$ STD, from the first 200 tokens within every instance. The color green represents AUCs higher than 60.
@@ 6.1 Assumptions to Evaluate
As shown in Table 1, prior research has verified that 4 out of 8 assumptions can fail under certain conditions. Meanwhile, some assumptions have specific requirements, and their applicability depends on how well these requirements are met. Therefore, we focus on two unverified assumptions that have no such requirements for evaluation, limiting confounding factors and deferring the testing of other assumptions to future studies. We also validate one verified assumption (Assumption A1) to confirm the consistency of our findings with prior research.
Assumption A1: Absolute Probability. In the assumption that seen instances will have a lower perplexity and fewer low-likely (outlier) tokens, we measure perplexity by an instance’s first $k$ tokens (PPL_$k$) Carlini et al. (2021); Min $p$Token by the average token probabilities among $p\%$ least likely tokens Shi et al. (2023).
Assumption A4: Verbatim Memorization. We expand this assumption from the instance level to the token level, assuming an LM will memorize some tokens in seen instances. We measure the percentage of tokens in an instance ranked as the $k$ most likely in casual language modeling (Mem $k$). The $k$ value of 1 represents greedy decoding, and larger than 1 simulates the decoding with top $k$ token sampling.
Assumption A6: Generation Variation. We evaluate the assumption that, given a seen prefix sequence, the LM exhibits less variation (i.e., higher certainty) in predicting the next token under different token sampling strategies. Since lower entropy indicates greater certainty, we measure entropy over the top $k$ most likely tokens (Entropy $k$) (see Appendix A.4 for details).
@@ 6.2 Experiment Design
To enhance the generalizability of our results, we evaluate these assumptions using four different types of LLMs, with datasets used in different training phases, shown in Table 2. We also investigate the impact of model size on MIA performance, using seven Pythia models with parameter sizes ranging from 70M to 13B.
Except for the UltraChat dataset, each dataset consists of multiple smaller subsets from different domains. For our experiments, we randomly sample 9 subsets from each, and consider each subset as an individual dataset. To minimize distribution shifts between seen and unseen datasets, we randomly select 1,000 instances from the training split (seen) and 1,000 instances from the test split (unseen) within each dataset. If a test split is unavailable, we sample from the validation split. If there are fewer than 1,000 unseen instances, we use the entire test split, ensuring that each split contains at least 100 instances.
Following prior work Shi et al. (2023), we evaluate MIA performance using the area under the ROC curve (AUC) at the instance level, representing the probability that a seen instance has a better score (higher or lower) than an unseen instance Google .
@@ 6.3 Results
6.3.1 Within-Domain MIA
Table 3 shows the AUC for each MIA method across representative subsets (domains) of each dataset. Complete results for all datasets and domains are available in Appendix A.5.1.
On pretraining datasets, all metrics perform close to random guessing, with AUC close to 50. We also observed the same pattern with different sizes of Pythia LMs. Our results for Assumption A1 are consistent with critiques they received (see Section 4.2.1). We suspect that during pretraining, LMs are more likely to learn underlying data distributions, instead of memorizing specific instances.
On fine-tuning datasets, we observed a great variation in MIA AUC across domains. The best-performing metric, PPL_200 on the RE-2012temp dataset, can have an AUC as high as 99.4. This suggests that data contamination from memorizing training instances remains a risk during the fine-tuning phase. Overall, the performance of the perplexity-based metric improves as the number of tokens increases. This trend is linked to the fine-tuning process, where tokens at the beginning of the training instance serve as input prompts but are not explicitly learned during training.
6.3.2 Cross-Domains MIA with Data Distribution Shifts
Within the same domain, the similar average PPL between seen and unseen instances indicates that they have similar underlying distributions, but also a high variation (STD). However, PPL differs a lot across domains. We therefore examine the impact of distribution shifts from different domains on the MIA performance, with the scenario where seen and unseen instances are from different domains.
Figure 1: Average MIA AUC for the Pythia-6.9b model with PPL_200, when the seen and unseen instances are from different domains. The abbreviations represent the domains in Table 12 in the Appendix A.5.1.
We present the AUC with PPL_200 in Figure 1. The AUC is high in the top-right corner when seen instances are from a domain with lower average PPLs and unseen instances are from a domain with higher average PPLs. Conversely, the bottom-left corner has low AUCs. This indicates that the accuracy of PPL_200-based MIA highly depends on the domain difference, instead of the seen vs. unseen distinction. More information about the PPL_200 distribution within and across domains is in Appendix A.5.2. A similar trend is observed with other metrics (see Appendix A.5.2 & A.5.3).
7 Discussion
In our case study, we observed the evaluated MIA approaches perform well only on certain domains of datasets used during the fine-tuning phase, but not during the pretraining phase. This discrepancy may be attributed to the significantly larger dataset and batch sizes employed in the pretraining phase.
Together, our case study and prior research show that 6 out of the 8 assumptions listed in Table 1 can often be invalid under certain conditions. The other two unverified assumptions, key information memorization (A5) and answer change due to memorization (A8), depend on very specific requirements, complicating their evaluation. Overall, our findings suggest that MIA remains a challenging task.
The limited effectiveness of MIA in pretraining phases suggests detecting data contamination in benchmarks remains an important challenge for LLM evaluation. While poor MIA performance may indicate a lack of direct instance memorization, models could still learn underlying distributional patterns of benchmark data, enabling artificially high performance on the benchmark datasets Dekoninck et al. (2024). On the other hand, privacy and copyright risks persist, as LLMs might learn from sensitive/proprietary data (e.g., patented concepts) without triggering MIA alarms.
@@ 8 Conclusion
In this study, we present a comprehensive survey of 50 studies focused on data contamination detection and their underlying assumptions. Our theoretical analysis reveals that these assumptions may not apply consistently across different contexts.
Our case studies showed that 3 out of the 8 assumptions are not universally applicable across all training phases and dataset domains, especially for datasets used in the pretraining stage. Our cross-domain MIA experiments additionally show that many assumptions measure an LM’s goodness of fit, which is not necessarily the result of instance memorization due to data contamination. Thus, detecting data contamination remains challenging.
@@ 9 Limitations
In this study, we reviewed 50 papers on data contamination detection, but there may be additional relevant studies not captured by our collection methods. We primarily focus on data contamination detection approaches for English LMs. Other detection approaches for non-English LMs, data modalities beyond text, and other machine learning techniques may exist and could potentially be transferable to English LMs.
Several factors influence an LM’s ability to memorize an instance. The relationship between MIA performance and the training dataset (e.g., domain, size, batch size), as observed in our case study, may vary for other LMs and their respective datasets.
@@ 10 Ethical Considerations
In this work, we employed multiple LMs and their training sets, which may contain sensitive and/or identifiable information. For example, the Pile Gao et al. (2020) includes content crawled from the Internet. The BioMistral-NLU-7B and its training set contain datasets derived from de-identified clinical notes Fu et al. (2025). Therefore, we only downloaded the necessary instances and published the numerical results from our experiments. In our project GitHub, we only release our code for data sampling and MIA approaches to ensure reproducibility; we do not publish any actual data instances or LMs. We recommend the community to check the corresponding regulations before deploying the datasets and LMs for other purposes.
11 Acknowledgements
This work was supported by the National Institutes of Health (NIH)—National Cancer Institute (Grant Nos. 1R01CA248422-01A1), National Library of Medicine (Grant No. 2R15LM01320902A1), and National Center for Advancing Translational Sciences of the National Institutes of Health (Grant No. UL1TR002319). The content is solely the responsibility of the authors and does not necessarily represent the official views of the NIH.
We thank Chenxi Li, Tianmai Zhang, and the anonymous reviewers for their feedback during the paper revision.
@@ A.1 Table of Notations
$x$ A language instance, as a sequence of tokens.
@@
contamination, which takes two instances
$x$ and $x^{\prime}$ as inputs, and returns False (0) or
@@
two instances, $x$ and $x^{\prime}$ and outputs
a real value.
@@
contamination, as defined in Equation 1.
@@
least likely tokens in an instance $x$, based
on a given LM $M$.
$\tau$
@@
produced by $M$ under diverse, different
sampling strategies. given $x_{p}$
PPL_$k$
The perplexity from an instance’s first
$k$ tokens.
@@
The entropy of the top $k$ most likely tokens
for the next position, defined by Equation 10.
@@ A.2 Risks and Mitigation Approaches for Data Contamination
During our paper collection, we identify the relevant research on the risks and mitigation strategies for data contamination, which does not involve proposing or evaluating existing detection approaches for data contamination. While excluding those papers from the main text of this manuscript, we provide their citations in Table 5 in this Appendix.
@@
Haimes et al. (2024) Mitigating data contamination in benchmarks through retrospectively creating held-out datasets.
Kusa et al. (2024) Proposing an evaluation pipeline in Systematic Literature Review to mitigate data contamination.